Ticket Number: HEA-NOC/20020910-01 Ticket Status: CLOSE
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: Dublin PoP
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 20020918 0800 UTC+1
Ticket Closed: 20020920 15:21:07 UTC Problem End: 20020918 1000 UTC+1
Repatching HEAnet server and office LANs and migrating to new
gateway. Plan to install a 48-port card in the new switch
(Calpurnia) and cutover to use it as gateway for the VLANs
during the next regular maintenance window (08:00 Wed 18th Sep).
The following will be unavailable for periods during the maintenance
window: NCI, HEAnet IPv6 gateway, and services hosted at the Dublin PoP
(e.g. www.heanet.ie, listserv.heanet.ie, mail.heanet.ie, ftp.heanet.ie,
keyhole.heanet.ie). Actual downtime is not expected to exceed a few
20020911 0835 cw all switchovers complete.
20020911 1035 reverting temporarily to altnertiove gateway, some servers may be uncontactable
20020913 1700 reload of core switch to carry out IOS upgrade. Outage of not more than 10 minutes within the 1 our window. Commencing at 0800 20020916
20020916 0853 CW Reload of switch/router complete. Upgrade of IOS was successful
20020918 1000 DW VLANs switched over, and NCI and Salinger (IPv6
gateway) have been physically moved to the new switch. A new security
policy is in place; if you have difficulty accessing a HEAnet service
please let us know. Ticket remains open for monitoring.
20020920 16:20 CW all stable.
Time to Fix:
<1 minute per VLAN during 2 hour maintenance window. Additional Notes:
Will monitor for next 24 hours.